OpenStudy (ajanijones):

Last month, Edwin had $2325.36 that he divided into several bank accounts. He deposited into an account to save for college. He deposited into another account to save for a car. How much money did he deposit to save for college and a car altogether?

OpenStudy (ajanijones):

@midhun.madhu1987

OpenStudy (ajanijones):

Please help I cant seem to getn the answer or know how to do it.

OpenStudy (anonymous):

there are missing numbers or fractions

OpenStudy (anonymous):

pls include the entire question for me to help you

OpenStudy (ajanijones):

Ok 1 sec. first one is 1/4 second is 1/3

OpenStudy (ajanijones):

Are you there.?

OpenStudy (texaschic101):

1/4 ( 2325.36) + 1/3 (2325.36) = total deposits OR 2325.36/4 + 2325.36/3 = total deposits

OpenStudy (ajanijones):

So.

OpenStudy (ajanijones):

I did that but not sure.

OpenStudy (texaschic101):

2325.35/4 + 2325.36/3 = 581.34 + 775.12 = 1356.46

OpenStudy (texaschic101):

oops...I wrote it wrong but figured it right.. 2325.36/4 + 2325.36/3 answer is : 1356.46

OpenStudy (ajanijones):

So is that the answer.

OpenStudy (ajanijones):

Tex.

OpenStudy (texaschic101):

yes...he deposited $581.34 for college and $775.12 for a car, totaling : $ 1356.46

OpenStudy (ajanijones):

Can you help me with one for to see if its right.

OpenStudy (texaschic101):

I can try

OpenStudy (ajanijones):

A recipe for potting soil that makes enough to fill 4 bags calls for 6-1/4 lb of compost. Kailee wants to modify the recipe so that it fills 10 bags. How many more pounds of compost will Kailee need for the modified recipe than the original recipe calls for?

OpenStudy (ajanijones):

A. 2-17/32 B. 9--3/8 C.7-13/16 D 15-5/8

OpenStudy (texaschic101):

A.) 2 - 17/32 = 64/32 - 17/32 = 47/32 = 1 15/32 B.) 9 - (-3/8) = 9 + 3/8 = 72/8 + 3/8 = 75/8 = 9 3/8 C.) 7 - 13/16 = 112/16 - 13/16 = 99/16 D.) 15 - 5/8 = 120/8 - 5/8 = 115/8 = 14 3/8

OpenStudy (ajanijones):

So my answer is B.

OpenStudy (texaschic101):

oh...whats the question ?

OpenStudy (ajanijones):

Omg if I could I would give you a thousand medalzz

OpenStudy (ajanijones):

A recipe for potting soil that makes enough to fill 4 bags calls for 6-1/4 lb of compost. Kailee wants to modify the recipe so that it fills 10 bags. How many more pounds of compost will Kailee need for the modified recipe than the original recipe calls for?

OpenStudy (ajanijones):

sorry I have 1 more.

OpenStudy (ajanijones):

Can you check this

OpenStudy (ajanijones):

Melanie’s monthly pay is $5553.60. She usually uses 1/3 of that amount to pay her household expenses, and 2/5 of that amount to pay for tuition. She deposits the rest into her savings account. This month, Melanie deposited twice the amount of money into her savings account than she normally deposits. How much did she deposit into her savings account this month? I got 8145.28

OpenStudy (texaschic101):

4/(6 1/4) = 10/x --- (4 bags to 6 1/4 lbs compost = 10 bags to x lbs compost) change mixed number to improper fraction 4/(25/4) = 10/x cross multiply 4x = (25/4)(10) 4x = 250/4 x = (250/4) / 4 x = 250/4 * 1/4 x = 250/16 x = 15 5/9 lbs are needed for 10 bags modified recipe - original recipe 250/16 - 25/4 = 250/16 - 100/16 = 150/16 = 9 3/8 yep...its B ==================== 5553.60/3 = 1851.20 (household expenses) 2/5(5553.60) = 11107.2/5 = 2221.44 (tuition) 5553.60 - (1851.20 + 2221.44) = 5553.60 - 4072.64 = 1480.96 (into savings account) deposited twice the amount of money.. 2(1480.96) = 2961.92 I got something totally different...

OpenStudy (ajanijones):

so whats your answer.

OpenStudy (texaschic101):

hold on...let me check again 1/3 + 2/5 = 5/15 + 6/15 = 11/15...leaving 4/15 for account (4/15) * 5553.6 = 22214.4/15 = 1480.96 1480.96 * 2 = 2961.92 I am getting the same answer doing it different ways....and that answer is $ 2961.92

OpenStudy (ajanijones):

thx.